OpenStudy (anonymous):

how many real number solutions are there in this quation -8x^2-8x-2=0

OpenStudy (anonymous):

Δ=0 so we have 1 double solution which is -0,5

OpenStudy (anonymous):

Umm what?

OpenStudy (anonymous):

I am 12 and i am doing this for a freind

OpenStudy (anonymous):

i am in 6th grade no mater what u say i will not get it

OpenStudy (anonymous):

so can i plz just get the answer

OpenStudy (anonymous):

ok then write this : Δ=64-4*16=0 x=-8/16=-1/2=-0.5

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Wait that is the answer right?

OpenStudy (anonymous):

yes write this and you'll be ok ,whoever has this homework will understand it's a formula don't worry :)